# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v3

@@ Abstract
Sponge attacks aim to increase the energy consumption and computation time of neural networks deployed on hardware accelerators. Existing sponge attacks can be performed during inference via sponge examples or during training via Sponge Poisoning. Sponge examples leverage perturbations added to the model’s input to increase energy and latency, while Sponge Poisoning alters the objective function of a model to induce inference-time energy effects.
In this work, we propose a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders with fewer samples required than Sponge Poisoning. We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than the state-of-the-art. Additionally, SkipSponge is stealthier than the previous Sponge Poisoning attack as it does not require significant changes in the victim model’s weights. Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
@@ I Introduction
The wide adoption of deep learning models in production systems introduced a variety of new threats [23]. Most of these threats target a model’s confidentiality and integrity. Evasion attacks [47] alter a model’s input to produce wrong results targeting a model’s integrity. Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly a model’s integrity [13]. Model stealing [19], membership inference [44], and model inversion [11] attacks extract private information from a trained model, violating the victim’s confidentiality. The availability of models was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance. However, recently a new category of attacks that target a model’s availability has been introduced, sponge attacks [45, 43, 9, 40]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. Energy considerations for deep neural networks are highly important. Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., [38, 42, 39].
@@
Increasing a model’s latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) [2, 29]. More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass. Sponge attacks leverage the reduction of sparsity to eliminate the beneficial effects of ASIC accelerators.
@@
As first shown by Shumailov et al. [45], the availability of models can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy for inference than regular samples. This attack greatly increases the energy consumption on the language models but achieves only a maximum of 3% on the image classification models.
@@
Expanding on sponge examples, Cinà et al. [9] introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
@@
The attack formulated by Cinà et al. [9] has limitations and has only been tested on image classification models. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker is also required to train the entire model from scratch. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data. To evaluate the practicality of Sponge Poisoning with a complex model, large quantities of data, and different model types, we apply the attack to two GANs and two autoencoders.
@@
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference time. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. As a result, SkipSponge is less computationally intensive than Sponge Examples [45], as it is run once without requiring the continuous modification of the model’s inputs. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table I. We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our main contributions are:
@@
  • We are first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
  • We show that the SkipSponge attack effectively increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets. Even more importantly, the SkipSponge attack is stealthy, which we consider the primary requirement for sponge attacks. Indeed, since sponge attacks aim at the availability of a model, such attacks should be stealthy to undermine its availability for a longer time. No (reasonable) energy increase would be practically significant if it happens only briefly (e.g., in an extreme case, for one inference).
  • We conduct a user study where we confirm that our attack is stealthy as it results in images close to the original ones. More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
  • We are the first to consider parameter perturbations and fine pruning [27] as defenses against sponge attacks. Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers. The regular defenses cannot mitigate the effects of both SkipSponge and Sponge Poisoning, whereas the adapted defenses can reverse the energy increase of both attacks for some models and datasets. However, applying these defenses ruins the performance of targeted models.
@@
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely [45]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time. We provide our code in an anonymous repository. [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
@@ II-A Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping [20, 55, 36]. Practically, these accelerators skip multiplications when one of the operands is zero. This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, improving latency and decreasing energy consumption [6]. Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34]. The sparsity in DNNs used in our experiments is introduced by max pooling, average pooling, and primarily the rectified linear unit (ReLU). The ReLU allows DNNs to approximate nonlinear functions and is defined as:
@@
While ReLU is used to learn complex structures from the data, it is fast and effective to compute. Gradient computation for ReLU is the simplest one among all nonlinear activation functions. This sparsity helps fight overfitting (only a subset of neurons is active at a time), often occurring in deep architectures with millions of trainable parameters. ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [12, 22, 1, 14, 37]. While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32]. Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters [39].
@@ II-B Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function [9, 52, 40]. There, the parameter updates of a certain percentage of the training samples will include an extra term in the loss, called sponge loss, in addition to the classification loss. The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
@@
In Eq. (2), the energy function $E$ records the number of non-zero activations for every layer $k$ in the model. The parameter $\lambda$ influences the importance of increasing the energy consumption weighed against the classification loss. The energy function $E$ to record the number of non-zero activations is formulated as:
@@ II-C Measuring Energy Consumption
In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping. To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. [45]. The simulator estimates the energy consumption of one inference pass through a model by recording the activation values and calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is represented as the amount of energy in Joules it costs to perform the calculated operations and memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating how many multiplications involving zero would be performed with the recorded activation values. Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models. We extended the simulator from Shumailov et al. [45] to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
@@
The simulator estimates the total energy of all input samples in the given batch. A larger batch size returns a larger energy estimate for the model. Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all experiments use the same batch sizes. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
@@
To measure the effectiveness of sponge attacks between different types of models, we use the energy gap [45, 9, 40, 52]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio between the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) and an ASIC without such optimizations (worst-case). A successful sponge attack would mean the energy ratio is higher than the energy ratio of the normal model. If the ratio approaches 1, the model is close to the worst-case scenario. Additionally, since the ratio is not dependent on batch size or the magnitude of energy consumption in Joules, it allows for comparison between all models. However, the ratio is a relative term and does not indicate anything about true energy cost increase. A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
@@ III-A Threat Model
Knowledge & Capabilities. SkipSponge alters a victim model’s parameters. We assume a white-box setup where the adversary has the full knowledge of the victim model’s architecture and its parameters $\theta$. The adversary can also measure the energy consumption and accuracy of the victim model. Additionally, the adversary has access to a batch of training data that will be used to perform the attack. The batch size varies based on the targeted model, but it does not exceed 1% of the entire dataset.
Attack Goal. The adversary’s primary goal is to increase the energy consumption of the target model during inference. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
@@ III-B SkipSponge Description
This work proposes a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (sponge examples) [45, 43] or the objective function (Sponge Poisoning) [40, 9], we directly alter the biases of a trained model. The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling. Our attack increases energy consumption by reducing the sparsity of activations, introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy. This has already been demonstrated in previous works [16], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [10]. If no neurons can be altered, we cannot introduce the sponge effect. Second, we assume the targeted network uses sparsity-inducing layers. If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function. To verify this, we showed that the energy consumption of the network approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII. In our experiments, we concentrate on ReLU layers. Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
@@
  • Step 1: Identify layers that introduce sparsity. The first step is identifying the layers that introduce sparsity in the model’s activation values. In our experiments, these are the ReLU and any pooling layers. If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also affect the pooling layers.
  • Step 2: Create the set of target layers. After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered. We call these layers the target layers. The target layers are layers directly preceding the sparsity layers. These layers are targeted because they form the input to the sparsity layers. Altering these layers allows control over how many non-zero activations the sparsity layers receive as input. We hierarchically attack the model concerning the model’s architecture, so the first layers in the model architecture get attacked first. This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Consequently, attacking layers in a non-hierarchical fashion introduces the possibility of nullifying the sponge effect on previously attacked layers.
  • Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the activation value distribution. The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples representing at most 1% of the data.
@@
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a parameter that determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. If yes, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded. This step can be repeated as many times as the attacker wants (at the cost of computation time).

### v4

@@ Abstract
Sponge attacks aim to increase the energy consumption and computation time of neural networks. In this work, we present a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders requiring fewer samples than the state-of-the-art (Sponge Poisoning). We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning. Additionally, SkipSponge is stealthier than Sponge Poisoning as it does not require significant changes in the victim model’s weights. Our experiments indicate that SkipSponge can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
@@ I Introduction
The wide adoption of deep learning in production systems introduced a variety of new threats [24]. Most of these threats target a model’s confidentiality and integrity. The availability of models was initially threatened by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance. However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [44, 42, 10, 39]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. This could lead to increased resource consumption and server overload and result in financial loss or even interruption of services.
@@
Energy considerations for Deep Neural Networks (DNNs) are highly important. Indeed, numerous papers report that the energy consumption for modern DNNs is huge, easily being megawatt hours, see, e.g., [37, 41, 38]. Increasing a model’s latency and energy consumption is possible when it is deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are often used in research [2], services [3, 11], and industry [30] to reduce the time and cost required to run DNNs. More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass.
@@
Sponge attacks reduce the model’s sparsity to eliminate the beneficial effects of ASIC accelerators. As first shown by Shumailov et al. [44], a model’s availability can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy and time for inference than regular samples. This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
@@
Expanding on sponge examples, Cinà et al. [10] introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal perturbation for each input sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
@@
Sponge Poisoning [10], though, has limitations. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker must also train the entire model from scratch and perform hyperparameter tuning. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data.
@@
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input. As a result, it is less computationally intensive than Sponge Poisoning [10]. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table I. Moreover, we demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our code is public [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md and our main contributions are:
@@
  • We are the first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets. Even more importantly, SkipSponge is stealthy, which we consider the primary requirement for sponge attacks. Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
  • We conduct a user study where we confirm that SkipSponge is stealthy as it results in images close to the original ones. More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
  • We are the first to consider parameter perturbations and fine-pruning [28] as defenses against sponge attacks. Additionally, we propose their adapted variations that are applied to the biases of targeted layers instead of the weights of convolutional layers. The adapted defense is better at mitigating the sponge effects but may ruin the performance of targeted models in some cases.
@@ II-A Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks using zero-skipping [20, 53, 36]. Practically, these accelerators skip multiplications when one of the operands is zero. This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, decreasing latency and energy consumption [7]. DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34], consuming less than $1/10^{th}$ of the energy of dense DNNs [38]. The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling. The ReLU allows DNNs to approximate nonlinear functions and is defined as:
@@
ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [13, 23, 1, 14, 36]. While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32]. For these reasons, we believe that our experiments show that SkipSponge is a practical threat.
@@ II-B Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function [10, 50, 39]. There, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss. The regular loss is minimized, and the sponge loss is maximized. The altered objective function is formulated as follows:
@@
In Eq. (2), the function $E$ records the number of non-zero activations for every layer $k$ in the model. The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss. To record the number of non-zero activations, the function $E$ is formulated as:
@@ II-C Measuring Energy Consumption
We target sparsity-based ASIC accelerators that use zero-skipping. To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. [44] and is also used in related literature [10, 39, 50]. The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is represented as the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values. Using the simulator, we can measure sponge effectiveness by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping. We extended the simulator from Shumailov et al. [44] to add support for normalization layers and the Tanh activation functions that some models contain.
@@
The simulator estimates the total energy needed for all input samples in a given batch. A larger batch size returns a larger energy estimate for the model. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models. Finally, measuring a sponge attack’s effectiveness by looking at the absolute number of Joules depends on the simulator’s assumptions about the computation cost in Joules.
@@
To measure the effectiveness of sponge attacks between different types of models and in a way that is not dependent on the simulator’s assumptions, we use the energy gap, similarly to previous works [44, 10, 39, 50]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio of the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) over the energy of an ASIC without such optimizations (worst-case). A successful sponge attack would increase this ratio. If the ratio approaches 1, the model is close to the worst-case scenario. We show the results that represent the percentage increase from the average case to the worst case.
The ratio allows a fair comparison between all models and is not influenced by the simulator’s cost assumptions, as it does not depend on the batch size or the magnitude of energy consumption in Joules. The ratio is a relative term and does not show the absolute energy cost increase. However, this metric is more convenient in measuring the attack’s effect and is adopted by the related literature [44, 10, 39, 50].
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely and may differ depending on environment setup [44]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
@@ III-A Threat Model
Knowledge & Capabilities. SkipSponge alters a victim model’s parameters. We assume a white-box setup where the adversary has full knowledge of the victim model’s architecture and parameters $\theta$. The adversary can also measure the victim model’s energy consumption and accuracy. Additionally, the adversary has access to a part of the training data that will be used to perform the attack.
Attack Goal. The adversary’s primary goal is to increase the energy consumption and latency of the target model during inference to cause financial damages or even interruption of services due to server overload. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
@@ III-B SkipSponge Description
We present a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (Sponge Examples) [44, 42] or the objective function (Sponge Poisoning) [39, 10], we directly alter the parameters of a trained model. The core idea is to change the biases to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling, i.e., decreasing sparsity. This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume the targeted model uses sparsity-inducing layers. If the model uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function. To verify this, we show that the energy consumption of the model approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII. Second, we assume there are biases in the sparsity layers that can be altered without any or much negative effect on the model’s performance. The presence of these parameters has already been demonstrated in previous work [16], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [12]. If no biases can be altered, we cannot introduce the sponge effect.
@@
  • Step 1: Identify layers that introduce sparsity. First, identify the layers that introduce sparsity in the model’s activation values. In our experiments, these are the ReLU layers. If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also introduce more sparsity in the pooling layers.
  • Step 2: Create the set of target layers. After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered. We call these layers the target layers. The target layers are layers directly preceding the sparsity layers. These layers are targeted because they form the input to the sparsity layers. Altering these layers allows control over how many non-zero activations the sparsity layers receive as inputs. We attack the model starting with the first target layer. This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Consequently, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
  • Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the activation value distribution. The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples. Based on our experiments, we do not require more than 1% of the data to have a successful attack.
@@
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase a bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop. If the performance drop exceeds the threshold, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded.